Calculus I, Fall 2011 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 05

Assigned Oct 13, 2011.

1. Section 3.8: Read problem 24.

Also read page 190 about the limitation of Newton’s method. (i.e. When does it work and not work?)

2. Section 3.8: problems 11, 12.

3. Section 4.2: problems 31, 33, 57, 61, 63, 81.

4. Section 4.3: problems 17, 25, 34, 75, 78, 80.